### v1

@@ Abstract
Context:Stellar populations and their distribution differ widely across the Galaxy, which is likely to affect planet demographics. Our local neighbourhood is dominated by young, metal-rich stars in the galactic thin disc, while the stellar halo and galactic bulge host a large fraction of older, metal-poor stars.
Aims:We study the impact of these variations on planet populations in different regions of the Galaxy by combining a high-resolution galaxy formation simulation with state-of-the-art planet population synthesis models.
Methods:We construct a population model to estimate occurrence rates of different planet types, based on the New Generation Planet Population Synthesis by Emsenhuber et al., 2021. We apply this to a simulated Milky Way–Analogue in the HESTIA galaxy formation simulation. We study the planet occurrence rate in the metal-rich regions of the inner Galaxy, i.e. in the galactic bulge and thin disc, and contrast them to the frequencies in the more distant, metal-poor region like the thick disc and stellar halo.
Results:We find that the planet demographics in the metal-poor regions of the Milky Way–Analogue, differ strongly from the planet populations in the more distant, metal-poor regions. The occurrence rate of giant planets () is 10 to 20 times larger in the thin disc compared to the thick disc, driven by the low amounts of solid material available for planet formation around metal-poor stars. Similarly, low-mass Earth-like planets around Sun-like stars are most abundant in the thick disc, being 1.5 times more frequent than in the thin disc. Moreover, low-mass planets are expected to be abundant throughout the galaxy, from the central regions to the outer halo, due to their formation processes being less dependent on stellar metallicity. The planet populations differ more strongly around Sun-like stars compared to dwarfs with masses 0.3 – 0.5 , caused by a weaker correlation between [Fe/H] metallicity and planet mass.
Conclusions:
Populating The Milky Way:
@@ Conclusions:
The architectures of planetary systems are tightly connected with the properties of their host stars, including their effective temperature, age, and metallicity. This relationship is most pronounced in the established correlation between stellar metallicity [Fe/H] and the occurrence rates of giant planets. Multiple observational studies have shown that giant planets are more commonly found around metal-rich stars (Gonzalez, 1997; Fischer & Valenti, 2005; Johnson et al., 2010). Additionally, observations suggest a decline in the fraction of stars hosting planets and a decrease in planet multiplicity with increasing mass of the host star (Yang et al., 2020; He et al., 2021). Across the galaxy, stellar populations exhibit significant variations in both age and chemistry (Gilmore & Reid, 1983; Prochaska et al., 2000; Reddy et al., 2006; Schuster et al., 2006; Bensby et al., 2014; Masseron & Gilmore, 2015), implying that planet demographics could strongly depend on the local galactic environment.
As galaxies grow and evolve, their local chemical profiles undergo changes due to stellar evolution (McWilliam, 1997). Throughout their life cycles, stars enrich the interstellar medium with heavy elements. This process, along with the hierarchical growth of galaxies, results in a metallicity gradient across the galaxy (Lemasle et al., 2007; Luck & Lambert, 2011; Genovali et al., 2014), as well as the formation of distinct stellar populations (Gilmore & Reid, 1983; Bland-Hawthorn & Gerhard, 2016a). The Milky Way, for instance, is typically divided into four components: the galactic bulge, thin disc, thick disc, and halo, each characterized by unique stellar populations with varied chemical compositions. The halo is particularly known for containing some of the oldest and most metal-poor stars in the galaxy, with metallicities ranging from -7 ¡ [Fe/H] ¡ -0.5 (Beers & Christlieb, 2005; Frebel & Norris, 2015). In contrast, the thin disc hosts a younger stellar population with metallicities near solar levels, [Fe/H] 0 (Kilic et al., 2017). The thick disc, though not as metal-rich as the thin disc, is distinguished by its -enhanced stars (Wallerstein, 1962). The bulge, representing the most heterogeneous of these populations, spans a broad range of metallicities, including both extremely metal-poor and metal-rich stars (McWilliam, 2016).
@@
Planet formation is an intricate process that involves a wide range of physical phenomena, and considerable effort has been dedicated to modeling and simulation work. Key aspects of this process include the formation of planetesimals and planetary embryos, the evolution and eventual dissipation of the protoplanetary disc, and the growth of planets through mechanisms such as accretion, migration, and the development of planetary interiors and atmospheres (e.g., Pollack et al., 1996; Ida & Lin, 2004a, b; Bitsch et al., 2019; Johansen et al., 2019; Alessi & Pudritz, 2018; Alessi et al., 2020). To align these models with observational data, planet population synthesis has become a crucial methodology (Ida & Lin, 2004a; Mordasini et al., 2009a, b). This approach involves generating representative samples of synthetic exoplanets for specific stellar populations, which can then be systematically compared to the actual observed exoplanet data.
@@
The heterogeneity of stellar populations, and their connection to galaxy evolution, has been a topic of recent research, which has has benefited greatly from the advancements in simulation techniques (Crain & van de Voort, 2023). Hydrodynamical simulations of galaxy evolution, in particular, have had great success in reproducing the properties of galaxies in the local universe (e.g. Navarro & White, 1994; Vogelsberger et al., 2014; Schaye et al., 2015; Dubois et al., 2016), including Milky Way-like galaxies. Amongst the Milky Way-reproducing simulations, one successful approach has been to model the local (galaxy) environment through the use of constrained simulations. These simulations are able to generate Milky Way-Analogues within environments that closely resemble the Local Group (Hoffman, 2009; Gottloeber et al., 2010; Libeskind et al., 2020).
@@
In this work, we aim to combine the recent advances in both galactic modelling and planetary population synthesis. We employ the HESTIA suite of constrained hydrodynamical galaxy formation simulations to study planet populations throughout the Milky Way (Libeskind et al., 2020). HESTIA reproduces the Local Group, including the Milky Way, Andromeda (M31), and surrounding satellite galaxies within the correct cosmographic context. Being a zoom-in simulation, HESTIA creates a high-resolution Milky Way-like galaxy with a smoothing length of approximately 220 pc and a star particle mass resolution of . The simulated galaxy closely matches the Milky Way in terms of total mass, stellar mass, metallicity distribution, and size. We combine one of these simulated galaxies with the state-of-the-art New Generation Planet Population Synthesis (NGPPS) data, obtained from the Bern model for planet formation and evolution, which generates realistic planet populations for a wide range of host star masses and metallicities (Emsenhuber et al., 2021a), in order to study the demographics of exoplanet throughout the Milky Way.
@@
This paper is organized as follows: Section 2 details the Bern model and NGPPS planet population, including the construction of a planet assignment model to associate planet occurrence rates with the mass and metallicity of the host stars. Section 3 describes the simulated Milky Way–Analogue in HESTIA, comparing key properties with the actual Milky Way, and presenting a decomposition of the simulated galaxy into bulge, thin disc, thick disc, and halo. This section also details the integration of the planet assignment model from Section 2 into the Milky Way–Analogue. In Section 4, we present the resulting planet populations in different parts of the galaxy, focusing on Sun-like and low-mass stars. These results are discussed in Section 5, where we also highlighting limitations of our study. Finally, we summarise our findings in Section 6.
@@ 2 Bern Planet Model and Planet Populations
The New Generation Planet Population Synthesis (NGPPS) is based on the third generation of the Bern global model for planet formation and evolution, a semi-numerical approach to study the formation of planets and evolution of planetary systems (Emsenhuber et al., 2021a). The Bern model is based on the core-accretion paradigm (Pollack et al., 1996), and its third generation encompasses the growth of planets from (1 – 100) planet embryos with a mass of about into planets with masses ranging from Earth-like planets to super-Jupiter Giants. The planet embryos are immersed into fluid-like protoplanetary discs, which are described by gas and planetesimal fields from which the planets grow. These fields are described by continuous density fields, coupled to the planet embryos. In the current generation of the model, direct N-body interaction between the growing planets is also included, and Emsenhuber et al. (2021a) have found that embryo – embryo interactions increase the number of massive planets compared to isolated evolution.
@@
The model is initialised at the stage when planet embryos have formed; the growth stages from dust to pebbles and planetesimals are not included. The initial number and distribution of the embryos are therefore free parameters. The simulation starts with a formation phase in which the planet embryos interact with the gas and planetesimal fields. This stage lasts for a fixed period of 20Myr, after which the disc is considered dispersed. From this point on, the planets evolve isolated, affected by tidal migration and atmospheric escape. The assumed formation phase of 20Myr is longer than typical protoplanetary disc lifetimes and therefore encompasses most of the growth stage. However, the late-stage giant impact phase that can affect low-mass planets is not fully simulated, which can lead to an underestimation of the masses for rocky planets. Emsenhuber et al. (2021a) find that for a Minimum Mass Solar Nebula-like surface density of solids, planet formation is mostly complete within AU after 20Myr, while more distant planets have not yet reached their final state. The model provides a wealth of data about the planet populations in the systems, including their mass, distance to the host star, radii, luminosities and evolution tracks.
@@
A wide range of physical processes are included in the model, and correspondingly a large number of parameters need to be set. Some of these parameters are fixed to specific values (Emsenhuber et al., 2021b, see Table 1), however some parameters are considered variable and a suit of simulations has been run varying these values. Four quantities are varied randomly and dubbed Monte Carlo variables:
@@
Initial gas disc mass: The initial mass of the gas disc around the star, denoted , provides the building material for planet formation, with a more massive disc potentially leading to more numerous and larger planets. Emsenhuber et al. (2021b) opted for log-normal distributions (with median , and standard deviation ) based on Very Large Array (VLA) observations reported by Tychoniec et al. (2018), which focused on Class I discs in the Perseus star-forming region. They restrict the range of of masses to 0.004 – 0.16 , to guarantee that the discs have non-negligible mass and are self-gravitationally stable.
@@
External photo-evaporation rate: External photo-evaporation is the main parameter determining the lifetime of the protoplanetary disc. While the inner part of the disc is evaporated by the internal radiation of the host star, the outer regions (where most of the mass resides) is evaporated primarily by the Far-UV radiation of nearby massive stars (Matsuyama et al., 2003). Emsenhuber et al. (2021b) assume a fixed disc viscosity and a normal distribution (, ) for to reproduce observed protoplanetary lifetimes. With their parametrisation 50% of discs will have a lifetime Myr and virtually all discs are dispersed within 10Myr.
@@
Dust-to-gas ratio: The dust-to gas ratio determines the amount of solid material in the disc, which is crucial for the growth of solid planets and the cores of gas giants. In the Emsenhuber et al. (2021a) model, the dust-to-gas ratio is assumed to be a function of stellar metallicity given by , where [Fe/H] is the iron abundance and is the dust-to-gas ratio of the sun. Emsenhuber et al. (2021b) choose a normal distribution (, ) for the [Fe/H] distribution truncated to the range [-0.6, 0.5]. This encompasses the vast majority of stars in the solar neighbourhood. However, in the galactic center and stellar halo metal-poor stars exhibit significantly lower [Fe/H] abundances. This needs to be considered for the HESTIA planet population model.
@@
Inner edge of the gas disc: The position the inner edge of the protoplanetary (gas) disc is important for the final locations of close-in planets, since gas drag is a major driver of migration. Emsenhuber et al. (2021b) choose a log-normal distribution (, ) for this quantity.
@@
The mass of the planetesimal (i.e. solid material) disc is not a Monte Carlo variable itself, but of central importance to planet formation since it provides the building blocks for rocky planets and giant cores. Within the Bern model, it is given is fully characterised by the mass of the gas disc and the dust-to-gas ratio, . Furthermore, the embryos are initialised with a a mass of , and distributed uniformly in the logarithm of distance between and 40 AU.
Besides these quantities, two more parameters are varied systematically:
@@
Host star mass: The host star mass () is fixed to in the work by Emsenhuber et al. (2021b). However Burn et al. (2021) extend the analysis to low mass stars with 10%, 30%, 50% and 70% of the sun’s mass. The host star’s mass affects the planet evolution in direct ways, e.g. through radiation, but is also correlated with other relevant quantities, e.g. the mass of the protoplanetary disc. In the work by Burn et al. (2021), only a run is considered, the Monte Carlo variables follow the same distribution as the for the Sun-like star but re-scaled through some fixed prescriptions, with and . The metallicity and external photo-evaporation rate distributions are assumed to be identical to the solar-mass case. Stars more massive than the sun were not studied, due to their relative rarity and short lifespan.

### v2

@@ Abstract
Context. Stellar populations and their distribution differ widely across the Galaxy, which is likely to affect planet demographics. Our local neighbourhood is dominated by young, metal-rich stars in the galactic thin disc, while the stellar halo and galactic bulge host a large fraction of older, metal-poor stars.
Aims. We study the impact of these variations on planet populations in different regions of the Galaxy by combining a high-resolution galaxy formation simulation with state-of-the-art planet population synthesis models.
Methods. We constructed a population model to estimate occurrence rates of different planet types, based on the New Generation Planet Population Synthesis (NGPPS). We applied this model to a simulated Milky Way (MW) analogue in the HESTIA galaxy formation simulation. We studied the planet occurrence rate in the metal-rich regions of the inner Galaxy, namely, in the galactic bulge and thin disc. We compared these result with the frequencies in the more distant, metal-poor region such as the thick disc and stellar halo.
Results. We find that the planet demographics in the central, metal-rich regions of the MW analogue differ strongly from the planet populations in the more distant, metal-poor regions. The occurrence rate of giant planets () is 10 to 20 times larger in the thin disc compared to the thick disc, driven by the low amounts of solid material available for planet formation around metal-poor stars. Similarly, low-mass Earth-like planets around Sun-like stars are most abundant in the thick disc, being 1.5 times more frequent than in the thin disc. Moreover, low-mass planets are expected to be abundant throughout the galaxy, from the central regions to the outer halo, due to their formation processes being less dependent on stellar metallicity. The planet populations differ more strongly around Sun-like stars compared to dwarfs with masses 0.3 – 0.5 , caused by a weaker correlation between [Fe/H] metallicity and planet mass. However, it is important to note that the occurrence rates of low-mass planets are still uncertain, making our findings strongly model-dependent. Massive planets are more comprehensively understood and our findings are more robust. Nonetheless, other systematic effects have the potential to alter the giant planet population that we have not addressed in this study. We discuss some of these limitations and offer further directions for future research.
@@
The architectures of planetary systems are tightly connected with the properties of their host stars, including their effective temperature, age, and metallicity. This relationship is most pronounced in the established correlation between stellar metallicity [Fe/H] and the occurrence rates of giant planets. Multiple observational studies have shown that close-in giant planets are more commonly found around metal-rich stars (Gonzalez, 1997; Fischer & Valenti, 2005; Johnson et al., 2010). Additionally, observations suggest a decline in the fraction of stars hosting planets and a decrease in planet multiplicity in line with the increasing mass of the host star (Yang et al., 2020; He et al., 2021). Across the galaxy, stellar populations exhibit significant variations in both age and chemistry (Gilmore & Reid, 1983; Prochaska et al., 2000; Reddy et al., 2006; Schuster et al., 2006; Bensby et al., 2014; Masseron & Gilmore, 2015), implying that planet demographics could strongly depend on the local galactic environment.
As galaxies grow and evolve, their local chemical profiles undergo changes due to stellar evolution (McWilliam, 1997). Throughout their life cycles, stars enrich the interstellar medium (ISM) with heavy elements. This process, along with the hierarchical growth of galaxies, results in a metallicity gradient across the galaxy (Lemasle et al., 2007; Luck & Lambert, 2011; Genovali et al., 2014), as well as the formation of distinct stellar populations (Gilmore & Reid, 1983; Bland-Hawthorn & Gerhard, 2016a). The Milky Way (MW), for instance, is typically divided into four components: the galactic bulge, thin disc, thick disc, and halo, each characterized by unique stellar populations with varied chemical compositions. The halo is particularly known for containing some of the oldest and most metal-poor stars in the galaxy, with metallicities ranging from -7 ¡ [Fe/H] ¡ -0.5 (Beers & Christlieb, 2005; Frebel & Norris, 2015). In contrast, the thin disc hosts a younger stellar population with metallicities near solar levels, [Fe/H] 0 (Kilic et al., 2017). The thick disc, though not as metal-rich as the thin disc, is distinguished by its -enhanced stars (Wallerstein, 1962). The bulge, representing the most heterogeneous of these populations, spans a broad range of metallicities, including both extremely metal-poor and metal-rich stars (McWilliam, 2016).
@@
Planet formation is an intricate process that involves a wide range of physical phenomena, and considerable effort has been dedicated to modeling and simulation work. Key aspects of this process include the formation of planetesimals and planetary embryos, the evolution and eventual dissipation of the protoplanetary disc, and the growth of planets through mechanisms such as accretion, migration, and the development of planetary interiors and atmospheres (e.g. Pollack et al., 1996; Ida & Lin, 2004a, b; Bitsch et al., 2019; Johansen et al., 2019; Alessi & Pudritz, 2018; Alessi et al., 2020). To align these models with observational data, planet population synthesis has become a crucial methodology (Ida & Lin, 2004a; Mordasini et al., 2009a, b). This approach involves generating representative samples of synthetic exoplanets for specific stellar populations, which can then be systematically compared to the actual observed exoplanet data.
@@
The heterogeneity of stellar populations, and their connection to galaxy evolution is an area of recent research that has has benefitted greatly from the advancements in simulation techniques (Crain & van de Voort, 2023). Hydrodynamical simulations of galaxy evolution, in particular, have had great success in reproducing the properties of galaxies in the local universe (e.g. Navarro & White, 1994; Vogelsberger et al., 2014; Schaye et al., 2015; Dubois et al., 2016), including MW-like galaxies. Amongst the MW-reproducing simulations, one successful approach has been to model the local (galaxy) environment through the use of constrained simulations. These simulations are able to generate MW analogues within environments that closely resemble the Local Group (Hoffman, 2009; Gottloeber et al., 2010; Libeskind et al., 2020).
@@
In this work, we aim to combine the recent advances in both galactic modelling and planetary population synthesis. We employ the HESTIA suite of constrained hydrodynamical galaxy formation simulations to study planet populations throughout the MW (Libeskind et al., 2020). HESTIA reproduces the Local Group, including the MW, Andromeda (M31), and surrounding satellite galaxies within the correct cosmographic context. As a zoom-in simulation, HESTIA creates a high-resolution MW-like galaxy with a smoothing length of approximately 220 pc and a star particle mass resolution of . The simulated galaxy closely matches the MW in terms of total mass, stellar mass, metallicity distribution, and size. We combine one of these simulated galaxies with the state-of-the-art New Generation Planet Population Synthesis (NGPPS) data, obtained from the Bern model for planet formation and evolution, which generates realistic planet populations for a wide range of host star masses and metallicities (Emsenhuber et al., 2021a). Its goal is to study the demographics of exoplanet throughout the MW.
@@
This paper is organized as follows. Section 2 details the Bern model and NGPPS planet population, including the construction of a planet assignment model to associate planet occurrence rates with the mass and metallicity of the host stars. Section 3 describes the simulated MW analogue in HESTIA, comparing the key properties with those of the actual MW and presenting a decomposition of the simulated galaxy into bulge, thin disc, thick disc, and halo. This section also details the integration of the planet assignment model from Section 2 into the MW analogue. In Section 4, we present the resulting planet populations in different parts of the galaxy, focussing on Sun-like and low-mass stars. These results are discussed in Section 5, where we also highlight the limitations of our study. Finally, we summarise our findings in Section 6.
@@ 2 Bern planet model and planet populations
The third generation of the Bern global model for planet formation and evolution, NGPPS, is based on a semi-numerical approach to study the formation of planets and evolution of planetary systems (Emsenhuber et al., 2021a). The Bern model is based on the core-accretion paradigm (Pollack et al., 1996), which is thought to be the primary formation pathway for low-mass planets, and a majority of massive planets. Specifically, it does not include giant planet formation through gravitational instability (Boss, 1997), which is thought to be a potential origin for very wide orbits ( AU, Rafikov, 2005; Schib et al., 2021) and very massive planets (, Schlaufman, 2018). The third generation of the model encompasses the growth of planets from (1 – 100) planet embryos with a mass of about into planets with masses ranging from Earth-like planets to super-Jupiter giants. The planet embryos are immersed into fluid-like protoplanetary discs, which are described by gas and planetesimal fields from which the planets grow. These fields are described by continuous density fields, coupled to the planet embryos. In the current generation of the model, direct N-body interaction between the growing planets is also included, and Emsenhuber et al. (2021a) have found that embryo-embryo interactions increase the number of massive planets compared to isolated evolution.
@@
The model is initialised at the stage when planet embryos have formed; the growth stages from dust to pebbles and planetesimals are not included. The initial number and distribution of the embryos are therefore free parameters. The simulation starts with a formation phase in which the planet embryos interact with the gas and planetesimal fields. This stage lasts for a fixed period of 20Myr, after which the disc is considered dispersed. From this point on, the planets evolve isolated, affected by tidal migration and atmospheric escape. The assumed formation phase of 20Myr is longer than typical protoplanetary disc lifetimes and therefore encompasses most of the growth stage. However, the late-stage giant impact phase that can affect low-mass planets is not fully simulated, which can lead to an underestimation of the masses for rocky planets. Emsenhuber et al. (2021a) found that for a minimum-mass solar nebula-like surface density of solids, planet formation is mostly complete within AU after 20Myr, while more distant planets have not yet reached their final state. The model provides a wealth of data about the planet populations in the systems, including their mass, distance to the host star, radii, luminosities, and evolution tracks.
@@
A wide range of physical processes are included in the model and, correspondingly, a large number of parameters need to be set. Some of these parameters are fixed to specific values (Emsenhuber et al., 2021b, see Table 1); however, some parameters are considered to be variable and a suite of simulations has been run varying these values. Four quantities are varied randomly and dubbed Monte Carlo variables.
@@
Initial gas disc mass: The initial mass of the gas disc around the star, denoted as , provides the building material for planet formation, with a more massive disc potentially leading to more numerous and larger planets. Emsenhuber et al. (2021b) opted for log-normal distributions (with median , and standard deviation ) based on Very Large Array (VLA) observations reported by Tychoniec et al. (2018), which focussed on Class I discs in the Perseus star-forming region. They restricted the range of of masses to 0.004 – 0.16 to guarantee that the discs have non-negligible mass and are self-gravitationally stable.
@@
External photo-evaporation rate: The external photo-evaporation measurement is the main parameter determining the lifetime of the protoplanetary disc. While the inner part of the disc is evaporated by the internal radiation of the host star, in the outer regions (where most of the mass resides) evaporation is initiated primarily by the far-ultraviolet (FUV) radiation of nearby massive stars (Matsuyama et al., 2003). Emsenhuber et al. (2021b) assumed a fixed disc viscosity and a normal distribution (, ) for to reproduce observed protoplanetary lifetimes. With their parametrisation 50% of discs will have a lifetime of Myr and virtually all discs are dispersed within 10Myr.
@@
Dust-to-gas ratio: The dust-to gas ratio, , determines the amount of solid material in the disc, which is crucial for the growth of solid planets and the cores of gas giants. In the Emsenhuber et al. (2021a) model, the dust-to-gas ratio is assumed to be a function of stellar metallicity given by , where [Fe/H] is the iron abundance and is the dust-to-gas ratio of the sun. Emsenhuber et al. (2021b) chose a normal distribution (, ) for the [Fe/H] distribution truncated to the range [-0.6, 0.5]. This encompasses the vast majority of stars in the solar neighbourhood. However, in the galactic centre and stellar halo metal-poor stars exhibit significantly lower [Fe/H] abundances. This needs to be considered for the HESTIA planet population model.
@@
Inner edge of the gas disc: The position the inner edge of the protoplanetary (gas) disc, , is important for the final locations of close-in planets, since gas drag is a major driver of migration. Emsenhuber et al. (2021b) chose a log-normal distribution (, ) for this quantity.
@@
The mass of the planetesimal (i.e. solid material) disc, , is not a Monte Carlo variable itself, but of central importance to planet formation since it provides the building blocks for rocky planets and giant cores. Within the Bern model, this is fully characterised by the mass of the gas disc and the dust-to-gas ratio, . Furthermore, the embryos are initialised with a a mass of , and distributed uniformly in the logarithm of distance between and 40 AU. In addition to these quantities, two more parameters are varied systematically:
@@
Host star mass: The host star mass () is fixed to in the work by Emsenhuber et al. (2021b). However Burn et al. (2021) extend the analysis to low mass stars with 10%, 30%, 50% and 70% of the sun’s mass. The host star’s mass affects the planet evolution in direct ways, for instance, through radiation, but is also correlated with other relevant quantities, such as the mass of the protoplanetary disc. In the work by Burn et al. (2021), only a run is considered, the Monte Carlo variables follow the same distribution as the for the Sun-like star but re-scaled through some fixed prescriptions, with and . The metallicity and external photo-evaporation-rate distributions are assumed to be identical to the solar-mass case. Stars more massive than the sun were not studied, due to their relative rarity and short lifespan.